On February 1, 2021, Arrow Construction Company entered into a three-year construction contract to build a bridge for a price of $8,120,000. During 2021, costs of $2,040,000 were incurred with estimated costs of $4,040,000 yet to be incurred. Billings of $2,540,000 were sent, and cash collected was $2,290,000.

In 2022, costs incurred were $2,540,000 with remaining costs estimated to be $3,660,000. 2022 billings were $2,790,000 and $2,515,000 cash was collected. The project was completed in 2023 after additional costs of $3,840,000 were incurred. The company’s fiscal year-end is December 31. Arrow recognizes revenue over time according to percentage of completion.

Required:
1. Compute the amount of revenue and gross profit or loss to be recognized in 2021, 2022, and 2023 using the percentage of completion method.
2a. Prepare journal entries for 2021 to record the transactions described (credit "various accounts" for construction costs incurred).
2b. Prepare journal entries for 2022 to record the transactions described (credit "various accounts" for construction costs incurred).
3a. Prepare a partial balance sheet to show the presentation of the project as of December 31, 2021.
3b. Prepare a partial balance sheet to show the presentation of the project as of December 31, 2022.
Answer #1

Answer-1:

Calculation of Revenue & Gross Profit to be recognized:
(Percentage completion method)
2021 2022 2023 Note
Contract Price        8,120,000         8,120,000          8,120,000 A
Cost Incurred during the year        2,040,000         2,540,000          3,840,000 B
Estimated cost to complete        4,040,000         3,660,000                       -   C
Cost Incurred to date        2,040,000         4,580,000          8,420,000 D
(Sum total of the cost incurred during the year)
Estimated total cost        6,080,000         8,240,000          8,420,000 E=C+D
Percentage complete 33.55% 55.58% 100.00% F=D/E
(Cost Incurred to date/Estimated total cost)
Total revenue        2,724,474         4,513,301          8,120,000 G=A*F
Previously recognized revenue                     -           2,724,474          4,513,301 H
Revenue to be recognized        2,724,474         1,788,827          3,606,699 I=G-H
Gross Profit to be recognized           684,474           (751,173)            (233,301) J=I-B

Answer-2a & 2b:

2021 2022
Journal Entries Debit $ Credit $ Debit $ Credit $
Costs Incurred CIP (Work-in-progress)         2,040,000        2,540,000
Cash          2,040,000        2,540,000
Billing Accounts Receivable         2,540,000        2,790,000
Billing on CIP          2,540,000        2,790,000
Receiving Payments Cash         2,290,000        2,515,000
Accounts Receivable          2,290,000        2,515,000
Period end adjustments for Revenue Construction Expense         2,040,000
(Gross Profit entry) CIP (Work-in-progress)            684,474
Construction Revenue          2,724,474
Period end adjustments for Revenue Construction Expense        2,540,000
(Gross Profit entry) CIP (Work-in-progress)           751,173
Construction Revenue        1,788,827